OpenStudy (anonymous):

How do I solve 12 1/2 = 2pi/b?

OpenStudy (anonymous):

\(\Large{12 \frac{1}{2}=\frac{2 \pi}{b}\\ \frac{25}{2}=\frac{2 \pi}{b}}\) solving for b, cross multiply  can you solve it from here?